crosstheline0236:

A rectangular stained glass window is made up of identical triangular parts as pictured below. mc019-1.jpg What is the area of the shaded triangular section? 18 square inches 36 square inches 72 square inches 144 square inches

Allison:

I can't see the picture. @crosstheline0236

Allison:

48/3 = 16 36/4 = 9 Multiply those two, and divide by 2 because the rectangle is divided into 2 triangles. 16*9/2 = 144

Allison:

@crosstheline0236